A dehydrated patient needs a 6.8% saline IV. Unfortunately, the hospital only has bags of 4% and 7% saline solutions. How many liters of each of these solutions should be mixed together to yield 3 liters of the desired concentration?
Answer by stanbon(75887)

A dehydrated patient needs a 6.8% saline IV. Unfortunately, the hospital only has bags of 4% and 7% saline solutions. How many liters of each of these solutions should be mixed together to yield 3 liters of the desired concentration?
----------------------
Equation:
salt + salt = salt
0.04x + 0.07(3-x) = 0.068*3
-----
Multiply thru by 1000 and solve for "x":
40x + 70*3 - 70x = 68*3
2*3 = 30x
x = 0.2 liters (amt. of 4% solution needed)
---
3-x = 2.8 liters (amt. of 7% solution needed)
-------------------------------------------
Cheers,
Stan H.